OpenStudy (anonymous):

please help!! Task 2 Part 1. Create two radical equations: one that has an extraneous solution, and one that does not have an extraneous solution. Use the equation below as a model. a√x+b +c=d Use a constant in place of each variable a, b, c, and d. You can use positive and negative constants in your equation. Part 2. Show your work in solving the equation. Include the work to check your solution and show that your solution is extraneous. Part 3. Explain why the first equation has an extraneous solution and the second does not.

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

\[a\sqrt{x+b}+c=d\]

ganeshie8 (ganeshie8):

assign some numbers to a,b,c,d

OpenStudy (anonymous):

any number?

ganeshie8 (ganeshie8):

for part a, you want the equation to have extraneous solution so make \(d\) less than \(c\)

ganeshie8 (ganeshie8):

yes choose ur favorite numbers such that d is less than c

OpenStudy (anonymous):

2,4,7,6

ganeshie8 (ganeshie8):

just to make the calculations much simpler, how about picking 1,4,7,6 ?

OpenStudy (anonymous):

ok ya fine with me

ganeshie8 (ganeshie8):

\[1\sqrt{x+4}+7=6\] solve \(x\)

ganeshie8 (ganeshie8):

start by subtracting 7 both sides

OpenStudy (anonymous):

-5

ganeshie8 (ganeshie8):

doesn't look right, show ur work

OpenStudy (anonymous):

I used a calculator for it

ganeshie8 (ganeshie8):

\[1\sqrt{x+4}+7=6\] subtract 7 both sides, what do you get ?

ganeshie8 (ganeshie8):

you get : \[\sqrt{x+4}=-1\] right ?

OpenStudy (anonymous):

ya sorry my computer glitched

ganeshie8 (ganeshie8):

next square both sides so that radical disappears on left hand side : \[x+4=1\]

ganeshie8 (ganeshie8):

isolate \(x\)

OpenStudy (anonymous):

would it be 1? and this is for part one right?

ganeshie8 (ganeshie8):

how did you get 1 ?

OpenStudy (anonymous):

you said to square it.

ganeshie8 (ganeshie8):

Ohk, next isolate x

OpenStudy (anonymous):

how do u do that? sorry im not good at this

OpenStudy (anonymous):

-3

ganeshie8 (ganeshie8):

good, so x = -3 is what you get by solving check if it really satisfies the given radical equation

ganeshie8 (ganeshie8):

plugin x=-3 in the equation \[1\sqrt{x+4}+7=6\]

OpenStudy (anonymous):

ok so I just input -3 into the radical equation and solve right

ganeshie8 (ganeshie8):

just input x=-3 into the radical equation and see if you get same thing on both sides

OpenStudy (anonymous):

its not true its false

ganeshie8 (ganeshie8):

that means x=-3 is an extraneous solution

ganeshie8 (ganeshie8):

we're done with half of part a and part b

ganeshie8 (ganeshie8):

create another equation cleverly such that you dont get an extraneous solution

OpenStudy (anonymous):

ok so for part a we have one equation and it does not have an extraneous solution am I right? so I pick four more numbers?

ganeshie8 (ganeshie8):

the one we worked just now has an extraneous solution

ganeshie8 (ganeshie8):

yes pick four more numbers such that you DONT get an extraneous solution

OpenStudy (anonymous):

oh ok and that equation was when I input the -3 in it? im just trying to understand

ganeshie8 (ganeshie8):

for the second equation without extranesou solution, you may simply exchange c and d \[1\sqrt{x+4}+6=7\]

ganeshie8 (ganeshie8):

this equation should not give you extraneous solutions

ganeshie8 (ganeshie8):

start by subtracting 6 both sides

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

sorry I mean 1 @ganeshie8

ganeshie8 (ganeshie8):

yes! square both sides and solve x all the way

OpenStudy (anonymous):

I got -3

ganeshie8 (ganeshie8):

Yep!

ganeshie8 (ganeshie8):

plug that into the original equation and see if it satisfies

OpenStudy (anonymous):

its true

ganeshie8 (ganeshie8):

that means the solution is NOT extraneous

ganeshie8 (ganeshie8):

we're done with part a and part b

OpenStudy (anonymous):

ok just to get this straight the extraneous solution was